What is the mass percent of water in a hydrate of copper(II) sulfate that has (1.00x10^0) waters of hydration?
1
Expert's answer
2015-09-22T09:51:39-0400
W(H2O) = M(H2O)*100% / M(CuSO4*H2O) = 18*100%/178 = 10.11%